# Towards Fair In-Context Learning with Tabular Foundation Models

### v2

@@ Abstract
Tabular foundational models have exhibited strong in-context learning (ICL) capabilities on structured data, allowing them to make accurate predictions on test sets without parameter updates, using training examples as context. This emerging approach positions itself as a competitive alternative to traditional gradient-boosted tree methods. However, while biases in conventional machine learning models are well documented, it remains unclear how these biases manifest in tabular ICL. The paper investigates the fairness implications of tabular ICL and explores three preprocessing strategies—correlation removal, group-balanced demonstration selection, and uncertainty-based demonstration selection—to address bias. Comprehensive experiments indicate that uncertainty-based demonstration selection consistently enhances group fairness of in-context predictions. The source code for reproducing the results of this work can be found at https://github.com/patrikken/Fair-TabICL.
@@ 1 Introduction
Tabular data, represented in rows and columns, is a data modality widely used for prediction tasks in domains such as finance and healthcare (Asuncion et al., 2007). Tree-based models such as XGboost (Chen et al., 2015) and Gradient-Boosted Trees (Ke et al., 2017) have shown the strongest generalization performance on tabular data. Recently, with the emergence of foundation models, Deep Learning (DL) based models have challenged the dominance of tree-based models (Hollmann et al., 2025). Foundation models are transformer models (Vaswani et al., 2017) pretrained on massive amounts of data, learning various complex structures that enable in-context learning (ICL) with a few labelled data (Brown et al., 2020). In-context learning has mainly been used with large language models (LLMs) for natural language tasks, where for a given prediction task, the labelled samples are formatted into textual demonstration examples and provided as context information to a language model such as GTP-3 (Radford et al., 2019). Given the demonstrations as context, the language model can effectively predict the label of a test example without any model update or finetuning (Brown et al., 2020). Attempts have been made to perform ICL using LLMs on tabular data, where the rows of the table are serialized into text or sentences (Hegselmann et al., 2023). However, LLMs are not pretrained to fully capture the complex relations between rows and columns of tabular data. Even when fine-tuned on vast amounts of tabular data, LLM-based ICL is limited by the size of the context window (up to 32 or 64 shots). To overcome this, Hollmann et al. (2022) proposed a TabPFN, a transformer-based tabular foundation model fully pretrained on synthetic datasets. Followup versions of TabPFN—TabPFNv2 (Hollmann et al., 2025) and TabICL (Qu et al., 2025)— demonstrated competitive or better performance compared to the traditional tree-based model on a variety of tasks under ICL (Hollmann et al., 2025; Qu et al., 2025).
With their state-of-the-art performance and ICL capabilities, transformer-based tabular foundation models will likely be widely adopted for decision-making and trigger a shift in the learning paradigm. However, using ICL-based models in high-stakes decision-making settings requires a thorough investigation of potential negative social impact, e.g., discrimination based on sensitive attributes such as gender, race, or sexual orientation. It has been shown that traditional machine-learning models can perpetuate bias in the data (Mehrabi et al., 2022) even when demographic features are not used to train the model (Kenfack et al., 2024a). Recent studies have demonstrated that ICL can also provide biased predictions (Hu et al., 2024; Bhaila et al., 2024). However, these works focus on LLM-based ICL with tabular data serialized into text format, thereby inheriting the drawbacks of foundational language models on tabular data (Hollmann et al., 2025). Moreover, tabular data is the main modality found in the fairness literature (Le Quy et al., 2022), and studies on fair ICL should use foundation models that can achieve state-of-the-art performance for tabular datasets (Qu et al., 2025).
@@
This paper investigates the fairness of ICL prediction using transformer-based tabular foundation models. First, our study reveals, perhaps unsurprisingly, that while these models focus on improving prediction accuracy, they can also amplify bias. Motivated by recent studies on the sensitivity of ICL performance—in terms of fairness and accuracy— to demonstration selection, we aim to address the following research question: What in-context selection/transformation method can improve the fairness of ICL predictions? In the fairness literature, fairness-enhancing methods are generally grouped into three categories: pre-processing, in-processing, and post-processing (Mehrabi et al., 2022). Projecting these categories into the ICL paradigm, pre-processing methods perform demonstration transformation or selection before predicting in context (Hu et al., 2024). In-processing methods would fine-tune or retrain the foundation model with fairness constraints (Robertson et al., 2024). Post-processing methods would alter the ICL predictions to improve a given fairness metric (Hardt et al., 2016). Pre- and Post-processing methods are more computationally friendly since they do not require model updates. This motivates our choice to focus on the pre-processing techniques, and we leave post-processing interventions for future exploration. More specifically, we propose and investigate three pre-processing fairness interventions: (i) Correlation Remover, a method that alters each non-sensitive feature to reduce their correlation with the sensitive attribute; (ii) group-balanced111Underlined represent the method’s name throughout the paper and in the results. in-context selection, ensures that the in-context set is group-balanced; (iii) Uncertainty-based in-context selection, estimates the uncertainty of predicting the sensitive attribute of in-context samples and only selects samples with uncertain predictions. We provide intensive experiments on eight fairness benchmark datasets to investigate the effectiveness of each method in terms of fairness and accuracy. Our results reveal that the uncertainty-based method can provide better fairness performance across datasets, fairness metrics, and foundational models, with a marginal impact on accuracy. Our contribution can be summarized as follows:
@@
While most existing studies focus on fair LLM-based ICL with serialized tabular data, we provide, to our knowledge, the first investigation into preprocessing methods for fair prediction in ICL using transformer-based tabular foundation models.
@@
We propose and investigate three pre-processing intervention methods to enforce fair ICL. These methods aim to reduce the information about the sensitive attributes of in-context samples. We demonstrate that uncertainty-based in-context sample selection can significantly improve the fairness of ICL predictions with a slight drop in accuracy.
@@
We release the code to ease the reproduction of the results and help researchers and practitioners integrate the proposed methods.
We consider a classification task with the giving training data where is an input feature vector, is the corresponding class label, and the corresponding demographic group. The goal is to learn a classifier to accurately predict the target given a sample while being fair w.r.t. demographic information . Several metrics have been proposed to measure fairness at the group or individual levels (Dwork et al., 2012). In this work, we focus on group fairness notions, measuring the performance disparity across different demographic groups, i.e., demographic parity, equalized odds, and equal opportunity (Hardt et al., 2016). A detailed description of these fairness metrics can be found in Appendix B.1.
@@ 2 Related works
Numerous methods have been proposed to enforce group fairness in classical machine learning models (Mehrabi et al., 2022; Mbiazi et al., 2023). Model-agnostic methods generally consist of modifying or reweighting the input data (pre-processing) to reduce information about the sensitive attribute (Madras et al., 2018; Creager et al., 2019; Kenfack et al., 2023; Kamiran & Calders, 2012; Celis et al., 2020; Balunović et al., 2021; Feldman et al., 2015; Kenfack et al., 2021) or altering the prediction outcomes (post-processing) to satisfy the desired fairness outcome (Hardt et al., 2016; Petersen et al., 2021). Some other methods adjust the training loss (in-processing) to involve fairness constraints during the training (Agarwal et al., 2018; Zhang et al., 2018; Roh et al., 2020). In contrast to prior works, this work focuses on pre-processing fairness intervention to ensure fair ICL predictions where the downstream model is a foundation model that makes predictions without model updates. We also note that model-agnostic fairness methods, such as pre-processing and post-processing, can be applied to ICL as they do not require model update. Benchmarking their performance on tabular foundation models under the ICL paradigm remains an open research question.
@@
In-context learning with tabular foundation models presents a notable advantage over traditional machine learning approaches by enabling models to adapt dynamically to new data without the need for retraining (Hollmann et al., 2022; Qu et al., 2025; Hollmann et al., 2025). Conventional ML methods typically depend on predefined training datasets, meaning that any alteration in the data or task necessitates a time-consuming and resource-intensive retraining process. In contrast, tabular foundation models utilize in-context learning to execute tasks based on the specific context of the data provided at inference time. This allows these models to interpret and process new tabular data with minimal prior preparation, facilitating more flexible and efficient decision-making (Hollmann et al., 2022). The advantages of this approach are particularly apparent in scenarios where data distributions change over time or when models must quickly adjust to various data tasks without undergoing retraining. Thus, as in-context learning emerges as a powerful tool for real-time, adaptive predictions in complex and dynamic environments, assessing and mitigating biases in the prediction can make its use more socially acceptable.
Studies on fair ICL use large language models with tabular datasets serialized into text or sentences (Bhaila et al., 2024; Hu et al., 2024; Ma et al., 2023). Specifically, Hu et al. (2024) investigates different group sampling methods to improve fairness on ICL predictions and suggests that prioritizing minority group demonstrations can improve the fairness performance of LLMs. Similarly, Bhaila et al. (2024) propose a data augmentation method to mitigate bias and select demonstration examples that can improve the fairness of ICL inference with LLMs. However, LLMs are not specifically trained to handle tabular data, and the overall performance of LLM-based ICL still lags behind well-tuned numeric models such as tree-based models (Hollmann et al., 2025). To our knowledge, this work represents the first initiative that investigates pre-processing fairness intervention on tabular foundation models. Robertson et al. (2024) proposed FairPFN, an in-processing method to improve counterfactual fairness (Kusner et al., 2017) at inference time. The proposed method trains the foundation model with synthetic fair datasets to eliminate the causal dependence on the sensitive attributes. During the pre-training stage, fairness is enforced by predicting the class label of the fair dataset using the biased data as context information. However, this method is computationally expensive since it requires retraining the foundation model. In contrast, we focus on pre-processing methods that do not require model update and aim to achieve group fairness metric instead of counterfactual fairness.
@@ 3.1 In-context Samples Transformation
Correlation remover (Feldman et al., 2015; Bird et al., 2020) is a preprocessing method that reduces the correlation between the sensitive and non-sensitive attributes before fitting the model. More specifically, a linear transformation is applied to each non-sensitive feature to reduce its correlation with the sensitive feature. We use the correlation remover as a preprocessing step over the training (in-context example) and testing sets before performing in-context prediction. Ultimately, transforming input features to reduce their linear dependency on the sensitive feature can reduce the reliance on sensitive features in the downstream models. However, as we will see in the results, nonlinear and complex downstream models can still infer the nonlinear dependencies over the sensitive feature and provide unfair results. A more detailed description of correlation remover is provided in Appendix B.2.
@@ 3.2 In-context Samples Selection
Kenfack et al. (2024b) demonstrated that models trained without fairness constraints can have better fairness properties when the training data consists of samples with uncertain sensitive attributes. We hypothesize that the uncertainty of the sensitive attribute prediction can be a good measure to select demonstrations that improve the fairness in-context predictions. To validate this, we measure the uncertainty of predicting the sensitive attribute in the demonstration set and use samples with high uncertainty for in-context learning. We focus on conformal prediction (Shafer & Vovk, 2008; Vovk et al., 2005) as uncertainty measure since it provides strong theoretical guarantees for the coverage. Instead of returning a single label, a conformal predictor returns a prediction set containing the true label with a probability of at least , with being a user-defined coverage parameter of the conformal prediction (Angelopoulos et al., 2023). For example, setting ensures the prediction set contains the true sensitive attribute value with at least probability. Specifically, samples with prediction sets containing more than one value are uncertain. Intuitively, the coverage parameter controls the fairness-accuracy tradeoff, with meaning no fairness intervention where all the datapoints are used and meaning maximal fairness intervention where only uncertain samples are included in the in-context examples.
Since conformal prediction is model agnostic, we considered both classical methods, e.g., Logistic Regression (LR), and foundation models, e.g., TabPFN for training the sensitive attribute classifier to measure the prediction uncertainties. Note that the method could be applied using other uncertainty measures, such as Monte Carlo dropout and confidence interval (Kenfack et al., 2024b). We focus on conformal prediction due to its rigorous theoretical guarantees, and it does not require a hyperparameter to threshold the level from which a prediction is considered uncertain (Angelopoulos et al., 2023).

### v3

@@ Abstract
Transformer-based tabular foundation models have recently demonstrated promising in-context learning (ICL) performance on structured data, emerging as competitive alternatives to gradient-boosted trees. However, the fairness implications of this new paradigm remain largely unexplored. We present the first investigation of fairness in tabular ICL, evaluating three recently proposed foundation models—TabPFNv2, TabICL, and TabDPT—on multiple benchmark datasets. To mitigate biases, we explore three pre-processing fairness-enhancing methods: correlation removal (decorrelating input features from the sensitive attribute), group-balanced sample selection (ensuring equal representation of protected groups in context examples), and uncertainty-based sample selection (prioritizing context examples with high sensitive-attribute prediction uncertainty). Our experiments show that the uncertainty-based strategy consistently improves group fairness metrics (e.g., demographic parity, equalized odds, and equal opportunity) with minimal impact on predictive accuracy. We release our code to facilitate reproducibility (https://github.com/patrikken/Fair-TabICL).
@@ 1 Introduction
Tabular data, represented in rows and columns, is a data modality widely used for prediction tasks in domains such as finance and healthcare (Asuncion et al., 2007). Tree-based models such as XGboost (Chen et al., 2015) and Gradient-Boosted Trees (Ke et al., 2017) have shown the strongest generalization performance on tabular data. Recently, with the emergence of foundation models, Deep Learning (DL) based models have challenged the dominance of tree-based models (Hollmann et al., 2025). Foundation models are models pretrained on vast datasets, without a specific task in mind, and they can be adapted across a wide range of downstream tasks. Large language models (LLMs) such as GPT-3 are common examples of foundation models, and they have demonstrated emerging capabilities such as in-context learning (ICL) with few labelled data (Brown et al., 2020). In-context learning (ICL) has primarily been applied to natural language tasks using large language models (LLMs). For example, in text classification, labeled examples are formatted as textual demonstrations and provided as context to a language model, enabling it to predict the label of a new instance without any parameter updates or fine-tuning (Radford et al., 2019; Brown et al., 2020). More recently, efforts have extended ICL to tabular data by serializing table rows into text or sentences (Hegselmann et al., 2023). However, since LLMs are not pretrained to model the complex structural relationships inherent in tabular data—such as interactions between rows and columns—their performance on large-scale tabular tasks still lags behind tree-based methods (Hegselmann et al., 2023).
Alternatively, recent work has proposed foundation models explicitly tailored for tabular data, achieving competitive performance with tree-based models while reducing the need for extensive model selection and hyperparameter tuning. For instance, TabPFN (Hollmann et al., 2022) is a transformer-based model pretrained on synthetic datasets, and its successor, TabPFNv2 (Hollmann et al., 2025), extends support to larger datasets with up to 10k samples. Similarly, TabICL (Qu et al., 2025), also pretrained on synthetic data, scales to datasets with up to 500k samples. These models leverage synthetic pretraining to encode a wide range of statistical priors, enabling effective target inference from in-context examples.
To better reflect the priors found in real-world datasets, other models incorporate real data during pretraining. Tabular Discriminative Pre-trained Transformer (TabDPT) (Ma et al., 2024) is pretrained directly on real-world datasets, while Real-TabPFN (Garg et al., 2025) builds on synthetic pretraining with additional fine-tuning on real data. These approaches generally yield improved performance by aligning the learned representations more closely with real-world data distributions.
Given their strong performance and in-context learning capabilities, transformer-based tabular foundation models will likely see widespread adoption in real-world decision-making tasks. This shift could mark a turning point in how tabular data problems are approached. However, the use of ICL-based models in high-stakes domains—such as healthcare, finance, or criminal justice—raises important ethical concerns. In particular, it is critical to assess their potential to perpetuate or even amplify existing social biases. Traditional machine learning models have already been shown to replicate biases present in the data (Mehrabi et al., 2022), and recent studies indicate that LLM-based ICL can also produce biased predictions (Hu et al., 2024; Bhaila et al., 2024). However, these studies rely on serialized representations of tabular data and therefore inherit the limitations of LLMs in handling tabular structures (Ma et al., 2024).
@@
This paper investigates the fairness of ICL prediction using transformer-based tabular foundation models. First, our study reveals, perhaps unsurprisingly, that while these models focus on improving prediction accuracy, they can also amplify bias. Motivated by recent studies on the sensitivity of ICL performance—in terms of fairness and accuracy— to demonstration selection, we aim to address the following research question: What in-context selection/transformation method can improve the fairness of ICL predictions? In the fairness literature, fairness-enhancing methods are generally grouped into three categories: pre-processing, in-processing, and post-processing (Mehrabi et al., 2022). Projecting these categories into the ICL paradigm, pre-processing methods perform demonstration transformation or selection before predicting in context (Hu et al., 2024). In-processing methods would fine-tune or retrain the foundation model with fairness constraints (Robertson et al., 2024). Post-processing methods would alter the ICL predictions to improve a given fairness metric (Hardt et al., 2016). Pre- and Post-processing methods are more computationally friendly since they do not require model updates. This motivates our choice to focus on the pre-processing techniques and leave post-processing interventions for future exploration. More specifically, we propose and investigate three pre-processing fairness interventions: (i) Correlation Remover (Feldman et al., 2015), a method that alters each input feature to reduce their correlation with the sensitive attribute; (ii) group-balanced111Underlined represent the method’s name throughout the paper and in the results. in-context selection, ensures that the in-context set is group-balanced; (iii) Uncertainty-based in-context selection, estimates the uncertainty of predicting the sensitive attribute of in-context samples and only selects samples with uncertain predictions. We performed intensive experiments on eight fairness benchmark datasets to investigate the effectiveness of each method in terms of fairness and accuracy. Our results reveal that the uncertainty-based method can provide better fairness performance across datasets, fairness metrics, and foundational models, with marginal impact on accuracy. Our contribution can be summarized as follows:
@@
While most existing studies focus on fair ICL with serialized tabular data, we provide, to our knowledge, the first investigation into preprocessing methods for fair prediction in ICL using transformer-based tabular foundation models.
@@
We propose and investigate three pre-processing intervention methods to enforce fair ICL predictions. These methods aim to reduce the information about the sensitive attributes of in-context samples. We demonstrate that uncertainty-based in-context sample selection can significantly improve the fairness of ICL predictions with a slight drop in accuracy.
@@
We release the code to ease reproduction of the results and help researchers and practitioners integrate the proposed methods.
@@ 2 Related works
Numerous methods have been developed to enforce group fairness in classical machine learning models (Mehrabi et al., 2022; Mbiazi et al., 2023; Kenfack et al., 2024a). These methods are often categorized as pre-processing, in-processing, or post-processing approaches. Model-agnostic methods in the pre-processing category typically modify or reweight the input data to reduce information correlated with sensitive attributes (Madras et al., 2018; Creager et al., 2019; Kamiran & Calders, 2012; Celis et al., 2020; Balunović et al., 2021; Kenfack et al., 2023). In contrast, post-processing techniques adjust the model’s prediction outcomes after training to satisfy fairness constraints (Hardt et al., 2016; Petersen et al., 2021). Finally, in-processing approaches embed fairness constraints directly into the training objective (Agarwal et al., 2018; Zhang et al., 2018; Roh et al., 2020). Unlike prior work, our approach focuses on pre-processing interventions applied in in-context learning (ICL) settings, where downstream predictions are made by foundation models without any model updates. We emphasize that model-agnostic methods—particularly pre- and post-processing—are especially suitable in ICL because they do not rely on access to or retraining of the model. However, their effectiveness in this setting, especially for tabular foundation models, remains largely unexplored and is the focus of our evaluation.
Tabular Fondation Models.
@@
In-context learning with tabular foundation models presents a notable advantage over traditional machine learning approaches by enabling models to adapt dynamically to new data without the need for retraining (Hollmann et al., 2022; Qu et al., 2025; Hollmann et al., 2025). Conventional ML methods typically depend on predefined training datasets, meaning that any alteration in the data or task necessitates a time-consuming and resource-intensive retraining process. In contrast, tabular foundation models utilize in-context learning to execute tasks based on the specific context of the data provided at inference time. This allows these models to interpret and process new tabular data with minimal prior preparation, facilitating more flexible and efficient decision-making (Hollmann et al., 2022). The advantages of this approach are particularly apparent in scenarios where data distributions change over time or when models must quickly adjust to various data tasks without undergoing retraining. Thus, as in-context learning emerges as a powerful tool for real-time, adaptive predictions in complex and dynamic environments, assessing and mitigating biases in the prediction can make its use more socially acceptable. Existing models are pre-trained using synthetic (Qu et al., 2025; Hollmann et al., 2022) or real-world data (Ma et al., 2024). Pre-training on real-world data often provides competitive or better performance, and the performance of synthetic pre-trained tabular foundation models can be boosted with continued pre-training on real-world data (Garg et al., 2025). In this work, we consider tabular foundation models pretrained on synthetic (Hollmann et al., 2025; Qu et al., 2025) and real-world data (Ma et al., 2024) and assess the fairness implications of each pretraining strategy.
Fairness in ICL.
Fairness in in-context learning has primarily been studied in the context of large language models (LLMs) applied to tabular data serialized as text(Bhaila et al., 2024; Hu et al., 2024; Ma et al., 2023). For instance, Hu et al. (2024) explore group-aware sampling strategies, finding that prioritizing minority group demonstrations improves fairness outcomes. Similarly, Bhaila et al. (2024) propose a data augmentation technique that guides demonstration selection to reduce bias during inference. While related in spirit, our approach differs in two key ways: (1) we focus on numerical tabular foundation models (rather than LLMs), and (2) our uncertainty-based selection method aims to reduce model reliance on sensitive attributes rather than optimize informativeness per se. Prior uncertainty-driven methods(Mavromatis et al., 2023; Kung et al., 2023) focus on selecting informative examples under a labeling budget, whereas we use uncertainty to guide fair demonstration selection. It is also important to highlight that LLMs, though flexible, are not optimized for tabular data and often perform worse than specialized numeric models, such as gradient-boosted trees (Hegselmann et al., 2023). As such, our work fills a key gap by investigating fairness interventions tailored to numeric tabular foundation models. To our knowledge, this is the first study to evaluate pre-processing fairness methods in this emerging model class.
Fairness-Aware Tabular Foundation Models.
Recent work by Robertson et al. (2024) introduced FairPFN, a TabPFN-style model trained to suppress the causal effect of sensitive attributes during pretraining. Their approach seeks counterfactual fairness (Kusner et al., 2017), ensuring that model predictions remain invariant when sensitive attributes are counterfactually changed. Our work is conceptually distinct in two respects. First, we do not require model retraining and rely on model-agnostic pre-processing methods, making our approach broadly applicable to any pretrained tabular foundation model. Second, we aim for group fairness, focusing on improving performance disparities between subgroups, rather than enforcing counterfactual invariance at the individual level.
@@ 3 Methodology
Problem Setup
We consider a classification task with the giving training data where is an input feature vector, is the corresponding class label, and the corresponding demographic group. The goal is to obtain a classifier , via ICL, to accurately predict the target given a sample while being fair w.r.t. demographic information . Several metrics have been proposed to measure fairness at the group or individual levels (Dwork et al., 2012). In this work, we focus on group fairness notions, measuring the performance disparity across different demographic groups, i.e., demographic parity, equalized odds, and equal opportunity (Hardt et al., 2016). A detailed description of these fairness metrics can be found in Appendix B.1.
@@ 3.1 In-context Samples Transformation
Correlation remover (Feldman et al., 2015; Bird et al., 2020) is a preprocessing method that reduces the correlation between the sensitive and non-sensitive attributes before fitting the model. More specifically, a linear transformation is applied to each non-sensitive feature to reduce its correlation with the sensitive feature. We use the correlation remover as a preprocessing step over the training (in-context example) and testing sets before performing in-context prediction. Ultimately, transforming input features to reduce their linear dependency on the sensitive feature can reduce the reliance on sensitive features in the downstream models. However, as we will see in the results, nonlinear and complex downstream models can still infer the nonlinear dependencies over the sensitive feature and provide unfair results. A more detailed description of correlation remover can be found in Appendix B.2.
@@ 3.2 In-context Samples Selection
Kenfack et al. (2024b) demonstrated that models trained without fairness constraints can have better fairness properties when the training data consists of samples with uncertain sensitive attributes. We hypothesize that the uncertainty of the sensitive attribute prediction can be a good measure to select demonstrations that improve the fairness of in-context predictions. To validate this, we measure the uncertainty of predicting the sensitive attribute in the demonstration set and use samples with high uncertainty for in-context learning. We focus on conformal prediction (Shafer & Vovk, 2008; Vovk et al., 2005) as uncertainty measure since it provides strong theoretical guarantees for the coverage. Instead of returning a single label, a conformal predictor returns a prediction set containing the true label with a probability of at least , with being a user-defined coverage parameter of the conformal prediction (Angelopoulos et al., 2023). For example, setting ensures the prediction set contains the true sensitive attribute value with at least probability. Specifically, samples with prediction sets containing more than one value are uncertain. Intuitively, the coverage parameter controls the fairness-accuracy tradeoff, with meaning no fairness intervention where all the datapoints are used and meaning maximal fairness intervention where only uncertain samples are included in the in-context examples. We show in Section 4.2.2 how the coverage parameter of the conformal predictor consistently controls the tradeoff between accuracy and fairness across fairness metrics and downstream foundational models.
Since conformal prediction is model agnostic, we considered both classical methods, e.g., Logistic Regression (LR), and foundation models, e.g., TabPFN for training the sensitive attribute classifier to measure the prediction uncertainties. Note that the method could be applied using other uncertainty measures, such as Monte Carlo dropout and confidence interval (Kenfack et al., 2024b). We focus on conformal prediction due to its rigorous theoretical guarantees, and it does not require a hyperparameter to threshold the level from which a prediction is considered uncertain (Angelopoulos et al., 2023). More details about uncertainty measurement with conformal prediction can be found in Appendix B.3.